# algorithmic

### v1

@@ 2.8.1 Conjugate updates (Cj)
Conjugate updates (Cj) provide a classical approach for computing the posterior by leveraging conjugate prior distributions. Conjugate updates occur when the functional form of the conditional prior $\tau({\bm{\theta}}_{t};\psi_{t},\mathcal{D}_{1:t-1})$ matches that of the measurement model $p({\bm{y}}_{t}\,|\,{\bm{\theta}},{\bm{x}}_{t})$ (Robert et al., 2007, Section 3.3). This property allows the posterior to remain within the same family as the prior, which leads to analytically tractable updates and facilitates efficient recursive estimation.
@@ 2.8.2 Linear-Gaussian approximation (LG)
The prior over model parameters is assumed to be:
$\tau_{t}({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t-1})=\mathcal{N}\left({\bm{\theta}}_{t}\,|\,\boldsymbol{\mu}_{t-1}^{(\psi_{t})},\boldsymbol{\Sigma}_{t-1}^{(\psi_{t})}\right),$ (21)
@@
$\lambda({\bm{\theta}}_{t};\psi_{t},\mathcal{D}_{1:t})$ $\propto\mathcal{N}({\bm{y}}_{t}\,|\,\bar{h}_{t}({\bm{\theta}}_{t},{\bm{x}}_{t}),\mathbf{R}_{t})\,\tau_{t}({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t-1})$ (23)
@@ 2.8.3 Variational Bayes (VB)
Variational Bayes (VB) is a popular method for approximating a posterior distribution of model parameters by choosing a parametric family (such as Gaussians) that is computationally tractable. The primary objective of VB is to minimise the Kullback-Leibler (KL) divergence between a candidate Gaussian distribution and the density $\lambda_{t}$. It can be shown that we can safely ignore the normalisation constant for $\lambda_{t}$, which is often computationally expensive, so we can replace $\lambda_{t}$ with its unnormalised form. We have the following optimisation problem for the posterior variational parameters:
$(\boldsymbol{\mu}_{t},\boldsymbol{\Sigma}_{t})=\operatorname*{arg\,min}_{\boldsymbol{\mu},\boldsymbol{\Sigma}}{\boldsymbol{\rm D}}_{\rm KL}\left(\mathcal{N}({\bm{\theta}}_{t}\,|\,\boldsymbol{\mu},\boldsymbol{\Sigma})\,\|\,p({\bm{y}}_{t}\,|\,{\bm{\theta}}_{t},{\bm{x}}_{t})\,\tau_{t}({\bm{\theta}}_{t};\,\psi_{t})\right),$ (25)
where $\tau_{t}({\bm{\theta}}_{t};\,\psi_{t})$ is the chosen prior distribution (M.3: prior).
An example of VB for neural network models is the Bayes-by-backpropagation method (BBB) of Blundell et al. (2015), which assumes a diagonal posterior covariance (more expressive forms are also possible). Nguyen et al. (2017) extended BBB to non-stationary settings. More recent approaches involve recursive estimation, such as the recursive variational Gaussian approximation (R-VGA) method of Lambert et al. (2022) which uses a full rank Gaussian variational approximation; the low-rank RVGA (L-RVGA) method of Lambert et al. (2023), which uses a diagonal plus low-rank (DLR) Gaussian variational approximation; the Bayesian online natural gradien (BONG) method of Jones et al. (2024), which combines the DLR approximation with EKF-style linearisation for additional speedups; the natural gradient Gaussian approximation (NANO) method of Cao et al. (2024), which uses a diagonal Gaussian approximation similar to VD-EKF in Chang et al. (2022); and the projection-based unification of last-layer and subspace estimation (PULSE) method of Cartea et al. (2023b), which targets different posterior densities for a subspace of the hidden layers and a full-rank covariance over the final layer of a neural network.
@@ 2.8.4 Alternative methods
$p({\bm{y}}_{t}\,|\,{\bm{x}}_{t},\psi_{t},\mathcal{D}_{1:t-1})=\int p({\bm{y}}_{t}\,|\,{\bm{\theta}}_{t},{\bm{x}}_{t})\,\tau({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t-1})\mathrm{d}{\bm{\theta}}_{t}.$ (27)
@@ 2.9.1 Discrete auxiliary variable (DA)
where $H:\mathbb{N}_{0}\to(0,1)$ is the hazard function. A popular choice is to take $H(r)=\pi\in(0,1)$ to be a fixed constant hyperparameter known as the hazard rate. The choice RL[inf]-PR is known as the Bayesian online changepoint detection model (BOCD).
@@
$p(\tilde{s}_{1:t}\,|\,s_{1:t-1},\mathcal{D}_{1:t-1})=\begin{cases}\pi&\text{if}\left([\tilde{s}_{1:t}\setminus\tilde{s}_{t}]=s_{1:t-1}\right)\text{ and }\tilde{s}_{t}=1,\\ 1-\pi&\text{if }\left([\tilde{s}_{1:t}\setminus\tilde{s}_{t}]=s_{1:t-1}\right)\text{ and }\tilde{s}_{t}=0,\\ 0&\text{otherwise},\end{cases}$ (29)
@@
which we use instead of $\nu_{t}(\psi_{t})$. For example, in RL above, ${\cal A}_{t-1}=\{r_{t-1}^{(k)}:k=1,\ldots,K\}$ are the unique $K$ most likely runlengths where the superscript represents the ranking according to $\nu_{t-1}(\cdot)$. Then, at time $t$, the augmented set $\bar{\cal A}_{t}$ becomes $({\cal A}_{t-1}+1)\cup\{0\}$, where the sum is element-wise, and we then compute the $K$ most likely elements of $\bar{\cal A}_{t}$ to define ${\cal A}_{t}$. In CPL, ${\cal A}_{t-1}$ contains the $K$ most likely sequences of changepoints, $\bar{\cal A}_{t}$ is defined as the collection of the $2\,K$ sequences where each sequence of ${\cal A}_{t-1}$ has a zero or one concatenated at the end. Finally, the $K$ most likely elements in $\bar{\cal A}_{t}$ define ${\cal A}_{t}$. This style of pruning is common in segmentation methods; see, e.g., Saatçi et al. (2010). However, other styles of pruning are also possible; see e.g., Li et al. (2021).
@@
A special case of the above is DA[1], where we employ a single hypothesis. In these scenarios, we set $\nu(\psi_{t})=1$ where $\psi_{t}$ is the most likely hypothesis. For example, RL[1] is a runlength $r_{t}$ with a single hypothesis. We provide another example next.
@@ 2.9.2 Continuous auxiliary variable (CA)
Here, we briefly discuss continuous auxiliary variables (CA). For some choices of $\psi_{t}$ and transition densities $p(\psi_{t}\,|\,\psi_{t-1},\mathcal{D}_{1:t-1})$, computation of (26) becomes infeasible. In these scenarios, $\nu_{t}$ is taken as a point-wise approximation of (26). We give an example below.
OU (OU rate with empirical Bayes estimate): for $\psi_{t}=\upsilon_{t}$, consider
@@
Because $\nu_{t}(\upsilon_{t})$ does not have closed-form solution and is continuous, the work of Titsias et al. (2024) takes
$\nu_{t}(\upsilon_{t})=\delta\left(\upsilon_{t}-\operatorname*{arg\,max}_{\upsilon\in[0,1]}p({\bm{y}}_{t}\,|\,{\bm{x}}_{t},\upsilon,\mathcal{D}_{1:t-1})\right)\,,$ (38)
where $\delta$ is the Dirac delta function. In practice, (38) is approximated by taking gradient steps towards the minimum. This is a form of empirical Bayes approximation, since we compute the most likely value of the prior $\psi_{t}$ after marginalizing out ${\bm{\theta}}_{t}$.
@@ 3 Unified view of examples in the literature
Table 2: List of methods ordered by publication date. The tasks are discussed in Section 2.3. We use the following abbreviations: SSSM means switching state space model; OCL means online continual learning; seg. means segmentation; preq. means prequential. Methods that consider two choices of (M.2: auxvar) are denoted by ‘X/Y’. This corresponds to a double expectation in (6) —one for each choice of auxiliary variable.
@@ 4 Experiments
Each experiment consists of a warmup period where the hyperparameters are chosen, and a deploy period where sequential predictions and updates are performed. In each experiment, we fix the choice of measurement model $h$ (M.1: likelihood) and posterior inference method (A.1: posterior), and then compare different methods with respect to their choice of (M.2: auxvar), (M.3: prior), and (A.2: weighting). For DA methods, we append the number of hypotheses in brackets to determine how many hypotheses are being considered. For example, RL[1]-PR denotes one hypothesis, RL[K]-PR denotes $K$ hypotheses, and RL[inf]-PR denotes all possible hypotheses. In all experiments, unless otherwise stated, we consider a single hypothesis for choices of DA. See Table 3 for the methods we compare.
@@
WoLF+RL-PR* (16) DA[inf] Special case of RL-PR with explicit choice of (A.1: posterior) which makes it robust to outliers. 4.3.2
@@
Table 3: List of methods we compare in our experiments. The first column, M.2–M.3, is defined by the choices of (M.2: auxvar) and (M.3: prior). The second column, Eq., references the equation that define M.2–M.3. The third column, A.2, determines the choice of (A.2: weighting). The fourth column, Description, provides a brief summary of the method. The fifth column, Sections, shows the sections where the method is evaluated. The choice of (M.1: likelihood) and (A.1: posterior) are defined on a per-experiment basis. (The only exception being WolF+RL-PR). For (M.2: auxvar) the acronyms are as follows: RL means runlength, CPP means changepoint probability, C means constant, and CPT means changepoint timestep. For (M.3: prior) the acronyms are as follows: PR means prior reset, OU means Ornstein–Uhlenbeck, LSSM means linear state-space model, Static means full Bayesian update, MMR means moment-matched prior reset, and OUPR means Ornstein–Uhlenbeck and prior reset. We use the convention in Hušková (1999) for the terminology abrupt/gradual changes.
@@ 4.1.1 Online regression for hour-ahead electricity forecasting
In this experiment, we consider the task of predicting the hour-ahead electricity load before and after the Covid pandemic. We use the dataset presented in Farrokhabadi et al. (2022), which has 31,912 observations; each observation contains 7 features ${\bm{x}}_{t}$ and a single target variable ${\bm{y}}_{t}$. The 7 features correspond to pressure (kPa), cloud cover (%), humidity (%), temperature (C) , wind direction (deg), and wind speed (KmH). The target variable is the hour-ahead electricity load (kW). To preprocess the data, we normalise the target variable ${\bm{y}}_{t}$ by subtracting an exponentially weighted moving average (EWMA) mean with a half-life of 20 hours, then dividing the resulting series by an EWMA standard deviation with the same half-life. To normalise the features ${\bm{x}}_{t}$, we divide each by a 20-hour half-life EWMA.
@@
For this experiment, we consider RL[1]-OUPR*, RL[1]-PR, C-ACI, and CPP-OU. For computational convenience, we plug in a point-estimate (MAP estimate) of the neural network parameters when making predictions in (6). For a fully Bayesian treatment of neural network predictions, see Immer et al. (2021); we leave the implementation of these approaches for future work.
@@
Figure 2: One day ahead electricity forecasting results for Fig. 1. The dotted black line corresponds to March 7 2020.

### v3

@@ 2.6.1 Conjugate updates (Cj)
Conjugate updates (Cj) provide a classical approach for computing the posterior by leveraging conjugate prior distributions. Conjugate updates occur when the functional form of the conditional prior $\pi({\bm{\theta}}_{t};\psi_{t},\mathcal{D}_{1:t-1})$ matches that of the measurement model $p({\bm{y}}_{t}\,|\,{\bm{\theta}},{\bm{x}}_{t})$ (Robert et al., 2007, Section 3.3). This property allows the posterior to remain within the same family as the prior, which leads to analytically tractable updates and facilitates efficient recursive estimation.
@@ 2.6.2 Linear-Gaussian approximation (LG)
The prior over model parameters is taken as:
$\pi({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t-1})=\mathcal{N}\left({\bm{\theta}}_{t}\,|\,\boldsymbol{\mu}_{t-1}^{(\psi_{t})},\boldsymbol{\Sigma}_{t-1}^{(\psi_{t})}\right),$ (23)
@@
$q({\bm{\theta}}_{t};\psi_{t},\mathcal{D}_{1:t})$ $\propto\mathcal{N}({\bm{y}}_{t}\,|\,\bar{h}_{t}({\bm{\theta}}_{t},{\bm{x}}_{t}),\mathbf{R}_{t})\,\pi({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t-1})$ (25)
@@ 2.6.3 Variational Bayes (VB)
Variational Bayes (VB) is a popular method for approximating a posterior distribution of model parameters by choosing a parametric family (such as Gaussians) that is computationally tractable. The primary objective of VB is to minimise the Kullback-Leibler (KL) divergence between a candidate Gaussian distribution and the density $q_{t}$. It can be shown that we can safely ignore the normalisation constant for $q_{t}$, which is often computationally expensive, so we can replace $q_{t}$ with its unnormalised form. We have the following optimisation problem for the posterior variational parameters:
$(\boldsymbol{\mu}_{t},\boldsymbol{\Sigma}_{t})=\operatorname*{arg\,min}_{\boldsymbol{\mu},\boldsymbol{\Sigma}}{\boldsymbol{\rm D}}_{\rm KL}\left(\mathcal{N}({\bm{\theta}}_{t}\,|\,\boldsymbol{\mu},\boldsymbol{\Sigma})\,\|\,p({\bm{y}}_{t}\,|\,{\bm{\theta}}_{t},{\bm{x}}_{t})\,\pi({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t-1})\right),$ (27)
where $\pi_{t}({\bm{\theta}}_{t};\,\psi_{t})$ is the chosen prior distribution (M.3: prior).
An example of VB for neural network models is the Bayes-by-backpropagation method (BBB) of Blundell et al. (2015), which assumes a diagonal posterior covariance (more expressive forms are also possible). Nguyen et al. (2017) extended BBB to non-stationary settings. More recent approaches involve recursive estimation, such as the recursive variational Gaussian approximation (R-VGA) method of Lambert et al. (2022) which uses a full rank Gaussian variational approximation; the low-rank RVGA (L-RVGA) method of Lambert et al. (2023), which uses a diagonal plus low-rank (DLR) Gaussian variational approximation; the Bayesian online natural gradient (BONG) method of Jones et al. (2024), which combines the DLR approximation with EKF-style linearisation for additional speedups; the natural gradient Gaussian approximation (NANO) method of Cao et al. (2024), which uses a diagonal Gaussian approximation similar to VD-EKF in Chang et al. (2022); and the projection-based unification of last-layer and subspace estimation (PULSE) method of Cartea et al. (2023b), which targets different posterior densities for a subspace of the hidden layers and a full-rank covariance over the final layer of a neural network.
@@ 2.6.4 Alternative methods
$p({\bm{y}}_{t}\,|\,{\bm{x}}_{t},\psi_{t},\mathcal{D}_{1:t-1})=\int p({\bm{y}}_{t}\,|\,{\bm{\theta}}_{t},{\bm{x}}_{t})\,\pi({\bm{\theta}}_{t};\,\psi_{t},\mathcal{D}_{1:t-1})\mathrm{d}{\bm{\theta}}_{t}.$ (29)
@@ 2.7.1 Discrete auxiliary variable (DA)
where $H:\mathbb{N}_{0}\to(0,1)$ is the hazard function. A popular choice is to take $H(r)=\kappa\in(0,1)$ to be a fixed constant hyperparameter known as the hazard rate. The choice RL[inf]-PR is known as the Bayesian online changepoint detection model (BOCD).
@@
$p(\tilde{s}_{1:t}\,|\,s_{1:t-1},\mathcal{D}_{1:t-1})=\begin{cases}\kappa&\text{if }\left([\tilde{s}_{1:t}\setminus\tilde{s}_{t}]=s_{1:t-1}\right)\text{ and }\tilde{s}_{t}=1,\\ 1-\kappa&\text{if }\left([\tilde{s}_{1:t}\setminus\tilde{s}_{t}]=s_{1:t-1}\right)\text{ and }\tilde{s}_{t}=0,\\ 0&\text{otherwise},\end{cases}$ (31)
@@
which we use instead of $\nu_{t}(\psi_{t})$. For example, in RL above, ${\cal A}_{t-1}=\{r_{t-1}^{(k)}:k=1,\ldots,K\}$ are the unique $K$ most likely runlengths where the superscript represents the ranking according to $\nu_{t-1}(\cdot)$. Then, at time $t$, the augmented set $\bar{\cal A}_{t}$ becomes $({\cal A}_{t-1}+1)\cup\{0\}$, where the sum is element-wise, and we then compute the $K$ most likely elements of $\bar{\cal A}_{t}$ to define ${\cal A}_{t}$. In CPL, ${\cal A}_{t-1}$ contains the $K$ most likely sequences of changepoints, $\bar{\cal A}_{t}$ is defined as the collection of the $2\,K$ sequences where each sequence of ${\cal A}_{t-1}$ has a zero or one concatenated at the end. Finally, the $K$ most likely elements in $\bar{\cal A}_{t}$ define ${\cal A}_{t}$. This style of pruning is common in segmentation methods; see, e.g., Saatçi et al. (2010), but other pruning are also possible, such as those proposed by Li et al. (2021), or sampling-based approaches; see e.g. Doucet et al. (2000).
@@
A special case of the above is DA[1], where we employ a single hypothesis. In these scenarios, we set $\nu(\psi_{t})=1$ where $\psi_{t}$ is the most likely hypothesis.
RL (Greedy runlength): For $\psi_{t}=r_{t}$ and DA[1], we take
$p(r_{t}\,|\,r_{t-1},\mathcal{D}_{1:t-1})=\begin{cases}1-\kappa&\text{if }r_{t}=r_{t-1}+1,\\ \kappa&\text{if }r_{t}=0,\\ 0&\text{otherwise}.\end{cases}$ (36)
Our choice of (A.2: weighting) is based on the marginal predictive likelihood ratio, which is derived from the computation of $p(r_{t}\,|\,\mathcal{D}_{1:t})$ under either either an increase in the runlength ($r_{t}^{(1)}=r_{t-1}+1)$ or a reset of the runlength ($r_{t}^{(0)}=0$). Under these assumptions, the form of $\nu_{t}(r_{t}^{(1)})$ is
$\nu_{t}(r_{t}^{(1)})=\frac{p({\bm{y}}_{t}\,|\,r_{t}^{(1)},{\bm{x}}_{t},\mathcal{D}_{1:t-1})\,(1-\kappa)}{p({\bm{y}}_{t}\,|\,r_{t}^{(0)},{\bm{x}}_{t},\mathcal{D}_{1:t-1})\,\kappa+p({\bm{y}}_{t}\,|\,r_{t}^{(1)},{\bm{x}}_{t},\mathcal{D}_{1:t-1})\,(1-\kappa)}.$ (37)
For details on the computation of (A.2: weighting), see Appendix A.3. For a detailed implementation of (M.2: auxvar) RL, (M.3: prior) OUPR, (A.2: weighting) DA[1], and (A.1: posterior) LG, see Algorithm 4 in the Appendix.
For example, RL[1] is a runlength $r_{t}$ with a single hypothesis. We provide another example next.
@@ 2.7.2 Continuous auxiliary variable (CA)
Here, we briefly discuss continuous auxiliary variables (CA). For some choices of $\psi_{t}$ and transition densities $p(\psi_{t}\,|\,\psi_{t-1},\mathcal{D}_{1:t-1})$, computation of (28) becomes infeasible. In these scenarios, we use simpler approximations. We give an example below.
CPP (Changepoint probability with empirical Bayes estimate): for $\psi_{t}=\upsilon_{t}$, consider
@@
The work of Titsias et al. (2024) takes $\nu_{t}(\upsilon_{t})=\delta(\upsilon_{t}-\upsilon_{t}^{*})$, where $\delta$ is the Dirac delta function and $\upsilon_{t}^{*}$ is a point estimate centred at the maximum of the marginal posterior predictive likelihood:
$\upsilon_{t}^{*}=\operatorname*{arg\,max}_{\upsilon\in[0,1]}p({\bm{y}}_{t}\,|\,{\bm{x}}_{t},\upsilon,\mathcal{D}_{1:t-1}).$ (42)
In practice, (42) is approximated by taking gradient steps towards the minimum. This is a form of empirical Bayes approximation, since we compute the most likely value of the prior after marginalizing out ${\bm{\theta}}_{t}$. The work of Galashov et al. (2024) considers a modified configuration with choice of (M.2: auxvar) ${\bm{\upsilon}}_{t}\in(0,1)^{m}$.
@@ 3 Unified view of examples in the literature
The methods presented in Table 3 can be directly applied to tackle any of the problems mentioned in Section 2.1. However, as choice of (M.1: model), we specify the task under which the configuration was introduced.444 In general, the components of BONE can be thought as the building blocks for new methods. Some of these combinations would not be useful, but they can be employed nonetheless.
@@
Table 3: List of methods ordered by publication date. The tasks are discussed in Section 2.1. We use the following abbreviations: SSSM means switching state space model; (O)CL means (online) continual learning; seg. means segmentation; preq. means prequential. Methods that consider two choices of (M.2: auxvar) are denoted by ‘X/Y’. This corresponds to a double expectation in (6)—one for each choice of auxiliary variable.
@@ 4 Experiments
Each experiment consists of a warmup period where the hyperparameters are chosen, and a deploy period where sequential predictions and updates are performed. In each experiment, we fix the choice of measurement model $h$ (M.1: model) and posterior inference method (A.1: posterior), and then compare different methods with respect to their choice of (M.2: auxvar), (M.3: prior), and (A.2: weighting). For DA methods, we append the number of hypotheses in brackets to determine how many hypotheses are being considered. For example, RL[1]-PR denotes one hypothesis, RL[K]-PR denotes $K$ hypotheses, and RL[inf]-PR denotes all possible hypotheses. In all experiments, unless otherwise stated, we consider a single hypothesis for choices of DA. See Table 4 for the methods we compare.
@@
WoLF+RL-PR* (18) DA[inf] Special case of RL-PR with explicit choice of (M.1: model) which makes it robust to outliers. 4.3.2
@@
Table 4: List of methods we compare in our experiments. The first column, M.2–M.3, is defined by the choices of (M.2: auxvar) and (M.3: prior). The second column, Eq., references the equation that define M.2–M.3. The third column, A.2, determines the choice of (A.2: weighting). The fourth column, Description, provides a brief summary of the method. The fifth column, Sections, shows the sections where the method is evaluated. The choice of (M.1: model) and (A.1: posterior) are defined on a per-experiment basis. (The only exception being WolF+RL-PR). For (M.2: auxvar) the acronyms are as follows: RL means runlength, CPP means changepoint probability, C means constant, and CPT means changepoint timestep. For (M.3: prior) the acronyms are as follows: PR means prior reset, OU means Ornstein–Uhlenbeck, LSSM means linear state-space model, Static means full Bayesian update, MMR means moment-matched prior reset, and OUPR means Ornstein–Uhlenbeck and prior reset. We use the convention in Hušková (1999) for the terminology abrupt/gradual changes.
@@ 4.1.1 Online regression for hour-ahead electricity forecasting
In this experiment, we consider the task of predicting the hour-ahead electricity load before and after the Covid pandemic. We use the dataset presented in Farrokhabadi et al. (2022), which has 31,912 observations; each observation contains 7 features ${\bm{x}}_{t}$ and a single target variable ${\bm{y}}_{t}$. The 7 features correspond to pressure (kPa), cloud cover (%), humidity (%), temperature (C) , wind direction (deg), and wind speed (KmH). The target variable is the hour-ahead electricity load (kW). To preprocess the data, we normalise the target variable ${\bm{y}}_{t}$ by subtracting an exponentially weighted moving average (EWMA) mean with a half-life of 20 hours, then dividing the resulting series by an EWMA standard deviation with the same half-life. To normalise the features ${\bm{x}}_{t}$, we divide each by a 20-hour half-life EWMA. The features are lagged by one hour.
@@
For this experiment, we consider RL[1]-OUPR* (our proposed method), RL[1]-PR (a classical method), C-ACI (a simple benchmark), and CPP-OU (a modern method). For computational convenience, we plug in a point-estimate (MAP estimate) of the neural network parameters when making predictions using $h$. More precisely, given $\psi_{t}$, we use $h({\bm{\theta}}_{t}^{*},{\bm{x}}_{t+1})$ to make a (conditional) prediction, where ${\bm{\theta}}_{t}^{*}=\operatorname*{arg\,max}_{\bm{\theta}}q({\bm{\theta}};\,\psi_{t},\mathcal{D}_{1:t})$. For a fully Bayesian treatment of neural network predictions, see Immer et al. (2021); we leave the implementation of these approaches for future work.
@@
Figure 4: One day ahead electricity forecasting results for Figure 3. The dotted black line corresponds to March 7 2020.